### v2

# Guardian: Detecting Robotic Planning and Execution Errors with Vision-Language Models

Abstract
Robust robotic manipulation requires reliable failure detection and recovery. Although current Vision-Language Models (VLMs) show promise, their accuracy and generalization are limited by the scarcity of failure data. To address this data gap, we propose an automatic robot failure synthesis approach that procedurally perturbs successful trajectories to generate diverse planning and execution failures. This method produces not only binary classification labels but also fine-grained failure categories and step-by-step reasoning traces in both simulation and the real world. With it, we construct three new failure detection benchmarks: RLBench-Fail, BridgeDataV2-Fail, and UR5-Fail, substantially expanding the diversity and scale of existing failure datasets. We then train Guardian, a VLM with multi-view images for detailed failure reasoning and detection. Guardian achieves state-of-the-art performance on both existing and newly introduced benchmarks. It also effectively improves task success rates when integrated into a state-of-the-art manipulation system in simulation and real robots, demonstrating the impact of our generated failure data. Code, Data, and Models available at https://www.di.ens.fr/willow/research/guardian/.
I Introduction
Despite recent advances in task planning and action execution enabled by Large Language Models (LLMs) [mistralsmall, grattafiori2024llama3herdmodels] and Vision-Language Models (VLMs) [zhu2025internvl3exploringadvancedtraining], current robotic manipulation systems remain vulnerable to a wide range of failures [sinha2023systemlevelviewoutofdistributiondata, Kawaharazuka_2024, kroemer2020reviewrobotlearningmanipulation]. They can suffer from incorrect task plans [huang2022language] due to LLM hallucinations [Huang_2025], perception errors such as confusing similar objects  [garcia2025generalizablevisionlanguageroboticmanipulation, goyal2023rvtroboticviewtransformer], and low-level control issues such as unstable grasps or slippage [wu2025robomind]. These failures often compound during task execution, severely hindering the reliability of robots in real-world scenarios. To improve robustness, there is a growing interest in models that can leverage vision and language to reason about task execution and detect robot failures [liu2023reflect, chen2024automatingrobotfailurerecovery, duan2025aha].
A key challenge in robotic failure detection is the lack of comprehensive failure datasets. Most real-world robot datasets consist of successful demonstrations [khazatsky2025droidlargescaleinthewildrobot, embodimentcollaboration2024openxembodimentroboticlearning, pumacay2024colosseumbenchmarkevaluatinggeneralization], offering limited insight into failure modes. Automatically collecting failures by running trained policies is slow and unsafe, while manually collecting failure data [chen2024automatingrobotfailurerecovery, bu2025agibot] is labor-intensive and often lacks diversity and realism. As a result, prior work [duan2025aha] largely relies on simulation to construct failure cases, which however suffers from the notorious sim-to-real gap [simtorealgapzhao2020]. Furthermore, they predominantly focus on execution errors, with limited coverage of planning failures [duan2025aha] and lack fine-grained failure reasons [liu2023reflect, agia2024unpackingfailuremodesgenerative, duan2025aha, chen2024automatingrobotfailurerecovery, bu2025agibot]. This scarcity of diverse, realistic, and richly annotated failure data significantly hampers the development and evaluation of failure detection methods.
Another challenge is effectively leveraging and reasoning over visual context for failure detection. REFLECT [liu2023reflect] converts images into textual descriptions and then leverages LLM reasoning to detect failures, but this multi-stage pipeline is vulnerable to error accumulation from incomplete or inaccurate image captions. Other recent approaches [agia2024unpackingfailuremodesgenerative, etukuru2024robotutilitymodelsgeneral] use VLMs to process images directly, yet typically rely on single-view observations, making them susceptible to occlusions and incomplete scene understanding. AHA [duan2025aha] addresses this by concatenating multi-view images across timesteps into a single grid-based representation; although this expands visual coverage, the compressed format can hinder fine-grained spatial-temporal understanding. Moreover, most existing methods output categorical labels without providing explicit reasoning traces, limiting interpretability and the ability to diagnose complex failure modes.
Figure 1: Illustration of our Guardian model - a VLM fine-tuned on our constructed failure datasets. It detects planning failures (top) and execution failures (bottom) in robotic manipulation.
To address these challenges, we introduce an automatic failure generation approach that produces diverse planning and execution failures, and we train a reasoning VLM Guardian on the resulting data for robust failure detection. Our method procedurally perturbs successful robot trajectories in both simulation and real-world settings to generate failures annotated with fine-grained categories and step-by-step reasoning traces. Using this pipeline, we construct three new datasets: RLBench-Fail (14K execution and 7K planning samples in the RLBench simulator [james2019rlbenchrobotlearningbenchmark]), BridgeDataV2-Fail (10K execution and 6K planning samples using the real-robot BridgeDataV2 data [walke2024bridgedatav2datasetrobot]), and UR5-Fail (570 execution and 370 planning samples), our collected real-robot failure dataset with policy-driven rollouts. We then develop Guardian, a reasoning VLM fine-tuned for failure detection. As illustrated in Fig. 1, Guardian formulates failure detection as a visual question-answering task, reasoning over task instructions, subtask descriptions, and the visual scene. It leverages high-resolution, multi-view images to provide fine-grained and interpretable failure categorization. Guardian achieves state-of-the-art performance on both RLBench-Fail and BridgeDataV2-Fail, and generalizes in a zero-shot manner to RoboFail [liu2023reflect] and our UR5-Fail dataset. We further demonstrate Guardian’s plug-and-play capability as a feedback module by integrating it into a robotic manipulation system 3D-LOTUS++ [garcia2025generalizablevisionlanguageroboticmanipulation]. Guardian significantly improves task success through planning and execution monitoring in both simulated and real-world tasks.
In summary, our contributions are three-fold:
  • We propose a novel robot failure generation approach that perturbs successful robot demonstrations in both simulation and real-world settings, and autonomously annotates failures with fine-grained categories and reasoning traces. This yields three new datasets for robot failure detection: RLBench-Fail, BridgeDataV2-Fail, and UR5-Fail.
  • We introduce Guardian, a fine-tuned VLM for planning and execution failure detection. Guardian integrates high-resolution, multi-view visual inputs with explicit reasoning to predict fine-grained categorization.
  • Guardian achieves state-of-the-art performance across four benchmarks, covering both in- and out-of-domain datasets. It further improves task completion of a vision-language manipulation framework in simulated and real robot tasks.
We will release datasets, code, and models.
II Related Work
Robot failure datasets. Existing robot failure datasets cover only limited settings. RoboFail [liu2023reflect] provides a small collection of hand-crafted simulation and real-world failures. Sentinel [agia2024unpackingfailuremodesgenerative] generates failures by rolling out a trained policy on out-of-distribution scenarios created through randomizing object scales and poses, but it includes only four tasks. AHA [duan2025aha] perturbs trajectories in RLBench simulator [james2019rlbenchrobotlearningbenchmark] to create over 49K image–query pairs, yet the data is restricted to simulation and does not include high-level planning failures. Moreover, none of these datasets provides step-by-step reasoning annotations, limiting the development of models capable of fine-grained failure reasoning. In this work, we propose an automated approach that generates large-scale datasets with diverse planning and execution failures, enriched with fine-grained categories and reasoning traces in both simulation and real-robot environments.
Failure detection in robotic manipulation. Traditional monitoring methods rely on explicit models of tasks, identifying model deviations [de1998execution, gianni2011unified]. Unlike LLM-based approaches, they depend on rigid model-based predictions. Recent works formulate failure detection as a Question Answering (QA) task, by zero-shot prompting LLMs and VLMs [agia2024unpackingfailuremodesgenerative, etukuru2024robotutilitymodelsgeneral, chen2024automatingrobotfailurerecovery, ma2024generative, shirai2024visionlanguageinterpreterrobottask, skreta2024replanroboticreplanningperception, mei2024replanvlmreplanningrobotictasks]. For instance, REFLECT [liu2023reflect] summarizes multi-sensory inputs as texts, including audio, and feeds the text to an LLM, but its reliance on consecutive off-the-shelf models leads to accumulated errors. Consequently, efforts have been made to fine-tune VLMs specifically for robotic failure detection. SuccessVQA [pmlr-v232-du23b] fine-tunes Flamingo [flamingo] using both simulated and real-world data, but evaluates only the final task success, while Guardian is designed for more granular monitoring, checking the execution of each subtask. Closest to our work, AHA [duan2025aha] fine-tunes LLaVA-1.5 [liu2023visualinstructiontuning] on multi-class failures generated in RLBench. AHA concatenates multiple camera viewpoints and keysteps into a single image, leading to compressed visual inputs. Unlike AHA and SuccessVQA, which treat failure detection as direct classification, Sentinel [agia2024unpackingfailuremodesgenerative] and Cosmos-Reason1 [nvidia2025cosmosreason1physicalcommonsense] frame it as a reason-first process with explicit chain-of-thought (CoT) [10.5555/3600270.3602070] reasoning but focus only on single-view binary failure detection. Sentinel proposes a zero-shot CoT prompting framework to leverage off-the-shelf VLMs, while Cosmos-Reason1 is fine-tuned for physical reasoning. Extending prior approaches, Guardian unifies planning and execution verification via chain-of-thought reasoning over separated multi-view inputs, for multi-class failure detection.
Learning-based robotic manipulation policies. Recent robotic policies have achieved remarkable performance on manipulation tasks [garcia2025generalizablevisionlanguageroboticmanipulation, embodimentcollaboration2024openxembodimentroboticlearning, ze20243ddiffusionpolicygeneralizable, black2025pi]. However, their real-world applicability is still limited by planning and execution errors leading to task failures [wu2025robomind]. Inspired by Manipulate-Anything [duan2024manipulateanythingautomatingrealworldrobots] and Robot Utility Models [etukuru2024robotutilitymodelsgeneral], which demonstrated significant benefits from integrating VLM verifiers into their pipelines, Guardian is designed as a plug-and-play verification module trained on specific failure data. It can seamlessly integrate into existing policies, enhancing robustness by verifying planning and execution stages.
III Robot Failure Datasets Construction
Figure 2: Failure Data Generation Pipeline. We introduce a novel generation pipeline generating failure cases both online in simulation (RLBench), and offline on the real-world dataset (BridgeDataV2). For each positive example, given its correct plan and successful trajectory, we generate a corresponding incorrect plan and unsuccessful trajectory.
III-A Data Sources
Simulated data enables controlled failure generation through procedural perturbations [duan2025aha], while real robot data reduces the sim-to-real gap but requires substantial human supervision [liu2023reflect]. To balance precise control and real-world fidelity, we use both simulated and real-robot datasets to construct robot failure datasets. We propose an automated method that derives planning and execution failures directly from successful demonstrations, avoiding manual failure collection. In both domains, tasks are decomposed into subtasks with corresponding video segments, which form the basis for generating failures. Fig. 2 (middle) illustrates successful episodes from the simulated and real robot datasets.
Simulated Data. We use the RLBench [james2019rlbenchrobotlearningbenchmark] simulator, selecting 52 tasks from RLBench-18Task [shridhar2022perceiveractormultitasktransformerrobotic] and GemBench [garcia2025generalizablevisionlanguageroboticmanipulation] benchmarks in our training data. For each task, we generate successful scripted trajectories with varied object placements and segment them into subtasks following 3D-LOTUS++ [garcia2025generalizablevisionlanguageroboticmanipulation].
Real Robot Data. We use BridgeDataV2 [walke2024bridgedatav2datasetrobot] with ECoT annotations  [zawalski2025roboticcontrolembodiedchainofthought], which provide fine-grained subtasks and object labels using large VLMs. We further clean these annotations automatically using heuristics and Mistral-Small-3.1-24B [mistralsmall] to filter episodes with missing targets or unreliable bounding boxes. To increase the number of successful trajectories, we augment data by reversing successful executions when applicable, by swapping their start and end images, and updating the associated instructions accordingly (e.g., “open drawer” becomes “close drawer”, “flip pot upright” becomes “flip pot upside down”). This yields approximately 20% additional successful demonstrations.
III-B Automated Failure Data Generation
We design failure modes based on established failure taxonomies [liu2023reflect, duan2025aha] and analysis of robot policy failures [wu2025robomind]. The failures are categorized into two types: planning and execution. A planning error denotes an incorrect decomposition of a task into subplans, whereas an execution error reflects unsuccessful completion of a subplan.
Planning Failures. As shown in Fig. 2 (top), we construct five types of planning failures:
  1. (1)
    Wrong object manipulated – some subtasks manipulate the wrong object.
  2. (2)
    Wrong object state or placement – some subtasks select the wrong target location, or state for the correct object.
  3. (3)
    Wrong order – one or several subtasks are not in the correct order, violating causal dependencies.
  4. (4)
    Missing subtask – required subtasks are missing from the plan, breaking task completeness.
  5. (5)
    Contradictory subtasks – some subtasks conflict with each other.
Types 1-3 are generated using an LLM (Mistral-Small-24B) to subtly alter the plan, while types 4-5 are created through rule-based perturbations. Each planning example comprises the task instruction, plan, and the initial front-view image.
Execution Failures. In simulation, we directly perturb subtask-level actions (Fig. 2, bottom left), leveraging the simulator’s precise control. A randomly selected subtask on the trajectory is modified using four failure modes:
  1. (1)
    No gripper close – the gripper is correctly positioned to grasp the object, but it fails to close its jaws.
  2. (2)
    Wrong object state or placement – the correct object is manipulated but ends in an incorrect state or placement.
  3. (3)
    Wrong object manipulated – the wrong object is used.
  4. (4)
    Imprecise grasping/pushing – the gripper attempts to grasp or push the correct object by moving toward it and closing its jaws, but misses due to inaccurate positioning.
Figure 3: Training examples. Guardian is trained on reasoning traces produced by our automatic data generation pipeline.
For real robot data, modifying actions directly is impractical due to current limitations of image editing and generation models. Therefore, we perturb the subtask text instruction paired with the pre-recorded trajectory segment (Fig. 2, bottom right) without direct robot control:
  1. (1)
    Task-execution semantic mismatch — an LLM (prompted with the original instruction and visible objects), or a rule-based preposition swap, generates a semantically altered instruction while preserving the start/end images.
  2. (2)
    Revert action — keep the instruction unchanged; replace the end image with the start one to show no progress.
Each execution example contains the task and subtask descriptions, plus pre-/post-action multi-view images.
III-C Chain-of-Thought (CoT) Generation
CoT reasoning has shown promise in improving the interpretability and performance of VLMs [zhang2024improve]. Therefore, we further explore whether reasoning can help failure detection. We introduce an automatic method to generate step-by-step CoTs for training reasoning models. For each sample, we first collect the object category, spatial location, and robot state from the RLBench simulator or from ECoT [zawalski2025roboticcontrolembodiedchainofthought] annotations, together with the corresponding failure reason. We then prompt a large reasoning-capable VLM (InternVL3-38B) [zhu2025internvl3exploringadvancedtraining] to generate step-by-step reasoning traces based on the initial text–image inputs and the aforementioned information. For planning samples, the model is instructed to sequentially verify each subtask and subsequently analyze the overall plan. For execution samples, the model is guided to describe the pre- and post-action images before assessing subtask completion. The reasoning trace contains 118 tokens on average. Fig. 3 illustrates training examples with chain-of-thoughts verifying plan correctness and subtask completion.
TABLE I: Robot failure datasets sizes across splits.
Training Validation Test Dataset Exec Plan Exec Plan Exec Plan RoboFail [liu2023reflect] - - - - 153 30 RLBench-Fail 12358 5808 1000 500 1000 500 BridgeDataV2-Fail 7830 4880 1000 500 1000 500 UR5-Fail 400 200 30 30 140 140
III-D Real-Robot, Policy-Driven Data Collection
We curate UR5-Fail, a real-robot dataset, collected using a UR5 arm with three cameras. We run the 3D-LOTUS++ policy [garcia2025generalizablevisionlanguageroboticmanipulation] on 34 tasks, recording initial and final multi-view images for each subtask. Subtasks are manually labeled as success or failure to obtain execution failure data. For planning failures, we annotate ground-truth plans and generate failures using the method described in Sec. III-B. Unlike RoboFail [liu2023reflect], which is single-view and relies solely on teleoperation, UR5-Fail is three-view and features autonomous policy rollouts yielding more realistic failures.
Figure 4: Failure mode distributions in real executions and our constructed data.
III-E Dataset Statistics and Evaluation
The resulting datasets, RLBench-Fail, BridgeDataV2-Fail, and UR5-Fail, contain balanced success/failure examples across both planning and execution, with fine-grained category labels and reasoning traces. Each dataset is split into training, validation, and test sets, with the validation and test sets featuring unseen tasks/environments to evaluate generalization. Dataset statistics are shown in Table I.
To measure the quality and diversity of our synthetic datasets, i.e., whether the generated failures reflect real policy execution, we run the 3D-LOTUS++ policy [garcia2025generalizablevisionlanguageroboticmanipulation] on 92 RLBench tasks and manually annotate failure modes for 3 failure episodes per task. As shown in Fig. 4, our designed failure modes reflect real failures, and the overall distribution of our synthetic and real failures remains similar.
IV Method: The Guardian Model
Figure 5: Left: Overview of the Guardian model architecture. Right: Integration of Guardian model into a robot manipulation pipeline for planning and execution verification.
IV-A Problem Formulation
We formulate robot failure detection as a visual question answering problem. For planning verification, given a high-level task instruction $T$, a proposed plan $P=(P_{1},\cdots,P_{N})$, and the initial visual context $I_{\text{start}}$, the model $\text{VLM}_{\text{plan}}$ must not only decide whether the plan is correct but also categorize the type of failure when it occurs:
$\text{VLM}_{\text{plan}}(I_{\text{start}},T,P)\rightarrow(B_{\text{plan}},C_{\text{plan}}),$ (1)
where $B_{\text{plan}}\in\{0,1\}$ indicates whether the plan is valid, and $C_{\text{plan}}$ specifies the failure category among the five planning failures constructed in Sec. III-B, shown in Fig. 4.
For execution verification, given the task goal $T$, a subtask description $P_{i}$, and the visual observations before and after execution, $I_{\text{start}}$ and $I_{\text{end}}$, the model $\text{VLM}_{\text{exec}}$ similarly outputs
$\text{VLM}_{\text{exec}}(I_{\text{start}},I_{\text{end}},T,P_{i})\rightarrow(B_{\text{exec}},C_{\text{exec}}),$ (2)
where $B_{\text{exec}}\in\{0,1\}$ indicates execution success, and $C_{\text{exec}}$ denotes the execution failure category among the five execution failures constructed in Sec. III-B, shown in Fig. 4.
IV-B Model Architecture
The Guardian model is built upon the state-of-the-art open-source VLM InternVL3-8B [zhu2025internvl3exploringadvancedtraining]. As illustrated in Fig. 5 (left), it consists of three key components: a text tokenizer that converts text prompts into discrete token embeddings, a visual encoder (InternViT-300M) that transforms individual images into visual embeddings, and a transformer-based LLM (Qwen2.5-7B) that processes the concatenated multimodal token sequence to predict the answer.
Rather than concatenating multiple images into a single grid-based image as in AHA [duan2025aha], Guardian processes each image independently through the visual encoder. This design preserves fine-grained spatial details within each image and allows the model to explicitly reason about spatial and temporal changes for more accurate failure detection. Furthermore, unlike SuccessVQA [nvidia2025cosmosreason1physicalcommonsense] and AHA [duan2025aha] that output direct classifications, Guardian leverages an explicit reasoning trace before concluding success or failure.
IV-C Model Training
We fine-tune Guardian on robot failure datasets using parameter-efficient Low-Rank Adaptation (LoRA) [hu2022lora], while freezing the visual encoder. Training minimizes cross-entropy loss for next-token prediction.
Although CoT has shown promise to improve performance, it brings additional computation overhead. Inspired by prior work [chen2025trainingstrategiesefficientembodied], we explore three strategies for incorporating CoT into failure detection: (1) Vanilla: a baseline model trained and evaluated to directly predict final answers (A) without CoT; (2) Thinking: the model is trained and inferred with explicit reasoning, always generating CoT before A; (3) Dropout: in training, the model alternates between generating CoT+A and directly predicting A, while at test time, only A is produced.
IV-D Integration into Robotic Manipulation Framework
Guardian can be seamlessly plugged into existing robotic manipulation pipelines as a verification layer without requiring any architectural modification. Without loss of generality, consider a modular robotic manipulation framework. As shown in Fig. 5 (right), Guardian can be inserted at each planning and subtask execution step to detect potential failures. Upon detection, it can trigger replanning or re-execute the corresponding motion policy to facilitate recovery.
V Experiments
V-A Experimental Setup
Evaluation datasets. We evaluate models on our newly constructed benchmarks RLBench-Fail, BridgeDataV2-Fail, and UR5-Fail, as well as the existing benchmark RoboFail [liu2023reflect], which is a manually created real-world failure dataset using a UR5 arm with a single camera.
Evaluation metric. We report the mean binary detection accuracy for each dataset (success/failure) in Sec. V-B and discuss multi-class categorization in Sec. V-C.
Implementation details. We trained our models using LoRA (rank 16, effective batch size 16) with the AdamW optimizer (weight decay 0.05), bf16 precision, and a cosine learning rate schedule peaking at $4\times 10^{-5}$. Training was conducted on the RLBench-Fail and BridgeDataV2-Fail sets (unless noted), completing in 5.5 hours on 4×H100 GPUs. During training, we randomly selected one or four views on RLBench-Fail data to mitigate overfitting to the multiple views. The best checkpoint is selected using the validation sets.
TABLE II: Comparison of SOTA failure detection models. Execution and Planning binary accuracies are reported. denotes numbers from the AHA paper. Guardian is fine-tuned from InternVL3-8B on our constructed datasets, boosting the performance of the base model by 23%.
(a) In-Domain datasets.
Model RLBench-Fail BDV2-Fail Exec Plan Exec Plan Large-scale generalist models Qwen3-VL-235B-A22B 0.59 0.83 0.75 0.86 GPT4.1 0.63 0.87 0.72 0.86 Specialized / small models CLIP+MLP [pmlr-v139-radford21a] 0.65 0.53 0.58 0.54 Sentinel [agia2024unpackingfailuremodesgenerative] 0.57 - 0.57 - Cosmos-Reason1-7B [nvidia2025cosmosreason1physicalcommonsense] 0.54 0.60 0.53 0.61 InternVL3-8B 0.59 0.70 0.66 0.71 Guardian-8B-Thinking 0.83 0.87 0.85 0.91
(b) Out-Of-Domain datasets.
Model RoboFail [liu2023reflect] UR5-Fail Exec Plan Exec Plan Large-scale generalist models Qwen3-VL-235B-A22B 0.82 0.70 0.79 0.84 GPT4.1 0.82 0.67 0.79 0.88 Specialized / small models CLIP+MLP [pmlr-v139-radford21a] 0.42 0.43 0.51 0.51 AHA-13B [duan2025aha] 0.64 - - - Sentinel [agia2024unpackingfailuremodesgenerative] 0.79 - 0.74 - Cosmos-Reason1-7B [nvidia2025cosmosreason1physicalcommonsense] 0.67 0.70 0.58 0.63 InternVL3-8B 0.77 0.53 0.71 0.75 Guardian-8B-Thinking 0.86 0.70 0.77 0.89
V-B Comparison with state of the art
Compared methods. We compare our approach against SOTA general and specialized vision–language baselines. General baselines include small-scale (InternVL3-8B [zhu2025internvl3exploringadvancedtraining]), and large-scale models (Qwen3-VL-235B-A22B-Thinking [qwen3vl2025], GPT4.1 [openai2025gpt41]). Specialized baselines comprise Cosmos-Reason1-7B [nvidia2025cosmosreason1physicalcommonsense], AHA-13B [duan2025aha], Sentinel [agia2024unpackingfailuremodesgenerative], and an MLP baseline using CLIP text-image features [pmlr-v139-radford21a]. We apply test-time CoT to the two large-scale models and Cosmos-Reason, as they are trained to produce reasoning traces and benefit from explicit reasoning at inference. Since AHA-13B is not publicly available, we report its binary accuracy on RoboFail from the original paper [duan2025aha]. Sentinel is limited to execution monitoring and is not evaluated for planning.
Results. Table II reports in-domain (ID) and out-of-domain (OOD) binary accuracies. Table II(a) shows that our Guardian model, fine-tuned from InternVL3-8B on failure datasets, surpasses/matches the larger Qwen3VL and GPT4.1 on in-domain execution and planning sets. In Table II(b) we evaluate models zero-shot on OOD datasets: Guardian attains 0.86/0.70 (Exec/Plan) binary accuracy on RoboFail and 0.77/0.89 on UR5-Fail, outperforming/matching larger general-purpose and specialized VLMs. Sentinel, applied atop Qwen3-VL, underperforms its base VLM, likely due to Sentinel’s rigid CoT prompt. Cosmos-Reason, trained on single-view binary failures, often overlooks fine-grained manipulator/object state changes. AHA, solely tuned to simulation failure data, does not generalize well to the real robot failures of RoboFail.
TABLE III: Impact of the Guardian fine-tuning data mix on the binary accuracy averaged over planning and execution.
Training Data RLBench BDV2 Robo UR5 RLBench BDV2 UR5 -Fail -Fail -Fail -Fail 0.65 0.69 0.65 0.73 0.80 0.72 0.69 0.74 0.66 0.87 0.72 0.69 0.85 0.88 0.78 0.83 0.85 0.88 0.74 0.85
Figure 6: Training data size impact on the Execution binary accuracy across benchmarks.
Figure 7: Confusion matrices for Guardian’s failure categorization for plans and executions. Values are percentages; rows denote ground-truth classes and columns the predicted classes.
V-C Failure Data Ablations
In this section, we examine the composition and scale of our data, and discuss the multi-class categorization.
Training data mixture. Table III shows the impact of fine-tuning datasets. Row 1 is the baseline InternVL 3-8B without fine-tuning. Fine-tuning on simulated RLBench-Fail data (row 2) significantly improves the performance on the same dataset, but the gains do not transfer well to the other real-world datasets due to the domain gap. A similar trend is observed when training solely on BridgeDataV2-Fail in row 3. In contrast, combining both datasets during fine-tuning in the fourth row leads to significant gains in domain and improves generalization to the OOD benchmarks. This indicates the importance of generating diverse planning and execution data. Further fine-tuning with the small real-robot UR5-Fail dataset improves performance on UR5-Fail while maintaining comparable accuracy on the other datasets.
Training data size. Fig. 6 presents the scaling behaviors of the training data size on the execution binary accuracies across RLBench-Fail, BridgeDataV2-Fail, RoboFail, and UR5-Fail. We observe a scaling behavior on in-domain and out-of-domain sets when fine-tuning on our procedurally generated data. This suggests that scaling the generated data size further could improve model performance.
Multi-class Categorization. Since our datasets contain fine-grained failure categorization, Guardian supports multi-class failure prediction. The confusion matrices in Fig. 7 show strong overall discrimination across categories. As shown in the Fig. 7(left), most execution categories are identified with high accuracy. The main weakness lies in the VLM’s visual grounding of subtle cues: no gripper close is confused with imprecise grasping/pushing, and wrong object state or placement is often mistaken for success. For planning shown in Fig. 7(right), wrong object state or placement, wrong order, missing subtask, and wrong object manipulated show higher confusion, particularly in long-horizon plans. Improving the model’s commonsense and temporal reasoning would further enhance its ability to capture these nuanced dependencies.
TABLE IV: Image representation impact on the Execution binary accuracy (number of views and image format). The AHA method [duan2025aha] equals row 3 (4-view concatenation).
Views
Image
Format
RLBench-Fail
Execution
1 concat 0.61
separated 0.63
4 concat (AHA [duan2025aha]) 0.72
separated 0.81
Figure 8: Verification with Guardian during online task execution on RLBench. Left: Successful correction of a generated plan. Right: Successful correction of a subtask execution.
V-D Failure Detection Method Ablations
In this section, we compare visual representation choices and reasoning strategies for VLM-based failure detection.
Multi-view images. Table IV investigates the image representation choice on failure detection. We fine-tune and evaluate InternVL3-8B solely on RLBench-Fail execution data. We vary the number of viewpoints (one or four) and the multi-image format, either separated as in Guardian, or concatenated into a single image as in AHA [duan2025aha]. For concatenated inputs, we represent the views as rows, and start/end images as columns. Using multi-view images consistently outperforms their single-view counterpart. In the single-view setting, separating or concatenating two images (start and end) yields comparable performance, as the resolution loss from concatenation is minimal. However, in the four-view setting, separating the images significantly boosts the performance by 9 points over concatenation. Concatenating images not only shifts from natural image distributions but also compresses visual information in limited image resolution.
TABLE V: Comparison of training and test-time strategies with and without CoT (Sec. IV-C). “A” denotes the final answer. We report the average binary accuracy and inference time [seconds/sample] on one H100.

### v3

@@
-Guardian: Detecting Robotic Planning and Execution Errors with Vision-Language Models
+Scaling Cross-Environment Failure Reasoning Data for Vision-Language Robotic Manipulation
 Abstract
 Robust robotic manipulation requires reliable failure detection and recovery.
-Although current Vision-Language Models (VLMs) show promise, their accuracy and generalization are limited by the scarcity of failure data.
-To address this data gap, we propose an automatic robot failure synthesis approach that procedurally perturbs successful trajectories to generate diverse planning and execution failures.
-This method produces not only binary classification labels but also fine-grained failure categories and step-by-step reasoning traces in both simulation and the real world.
-With it, we construct three new failure detection benchmarks: RLBench-Fail, BridgeDataV2-Fail, and UR5-Fail, substantially expanding the diversity and scale of existing failure datasets.
-We then train Guardian, a VLM with multi-view images for detailed failure reasoning and detection.
-Guardian achieves state-of-the-art performance on both existing and newly introduced benchmarks.
-It also effectively improves task success rates when integrated into a state-of-the-art manipulation system in simulation and real robots, demonstrating the impact of our generated failure data.
+Although recent Vision-Language Models (VLMs) show promise in robot failure detection, their generalization is severely limited by the scarcity and narrow coverage of failure data.
+To address this bottleneck, we propose an automatic framework for generating diverse robotic planning and execution failures across both simulated and real-world environments.
+Our approach perturbs successful manipulation trajectories to synthesize failures that reflect realistic failure distributions, and leverages VLMs to produce structured step-by-step reasoning traces.
+This yields FailCoT, a large-scale failure reasoning dataset built upon the RLBench simulator and the BridgeDataV2 real-robot dataset.
+Using FailCoT, we train Guardian, a multi-view reasoning VLM for unified planning and execution verification.
+Guardian achieves state-of-the-art performance on three unseen real-world benchmarks: RoboFail, RoboVQA, and our newly introduced UR5-Fail.
+When integrated with a state-of-the-art LLM-based manipulation policy, it consistently boosts task success rates in both simulation and real-world deployment.
+These results demonstrate that scaling high-quality failure reasoning data is critical for improving generalization in robotic failure detection.
 Code, Data, and Models available at https://www.di.ens.fr/willow/research/guardian/.
 I Introduction
-Despite recent advances in task planning and action execution enabled by Large Language Models (LLMs) [mistralsmall, grattafiori2024llama3herdmodels] and Vision-Language Models (VLMs) [zhu2025internvl3exploringadvancedtraining], current robotic manipulation systems remain vulnerable to a wide range of failures [sinha2023systemlevelviewoutofdistributiondata, Kawaharazuka_2024, kroemer2020reviewrobotlearningmanipulation].
-They can suffer from incorrect task plans [huang2022language] due to LLM hallucinations [Huang_2025], perception errors such as confusing similar objects  [garcia2025generalizablevisionlanguageroboticmanipulation, goyal2023rvtroboticviewtransformer], and low-level control issues such as unstable grasps or slippage [wu2025robomind].
-These failures often compound during task execution, severely hindering the reliability of robots in real-world scenarios.
-To improve robustness, there is a growing interest in models that can leverage vision and language to reason about task execution and detect robot failures [liu2023reflect, chen2024automatingrobotfailurerecovery, duan2025aha].
-A key challenge in robotic failure detection is the lack of comprehensive failure datasets.
-Most real-world robot datasets consist of successful demonstrations [khazatsky2025droidlargescaleinthewildrobot, embodimentcollaboration2024openxembodimentroboticlearning, pumacay2024colosseumbenchmarkevaluatinggeneralization], offering limited insight into failure modes.
-Automatically collecting failures by running trained policies is slow and unsafe, while manually collecting failure data [chen2024automatingrobotfailurerecovery, bu2025agibot] is labor-intensive and often lacks diversity and realism.
-As a result, prior work [duan2025aha] largely relies on simulation to construct failure cases, which however suffers from the notorious sim-to-real gap [simtorealgapzhao2020].
-Furthermore, they predominantly focus on execution errors, with limited coverage of planning failures [duan2025aha] and lack fine-grained failure reasons [liu2023reflect, agia2024unpackingfailuremodesgenerative, duan2025aha, chen2024automatingrobotfailurerecovery, bu2025agibot].
-This scarcity of diverse, realistic, and richly annotated failure data significantly hampers the development and evaluation of failure detection methods.
-Another challenge is effectively leveraging and reasoning over visual context for failure detection.
-REFLECT [liu2023reflect] converts images into textual descriptions and then leverages LLM reasoning to detect failures, but this multi-stage pipeline is vulnerable to error accumulation from incomplete or inaccurate image captions.
-Other recent approaches [agia2024unpackingfailuremodesgenerative, etukuru2024robotutilitymodelsgeneral] use VLMs to process images directly, yet typically rely on single-view observations, making them susceptible to occlusions and incomplete scene understanding.
-AHA [duan2025aha] addresses this by concatenating multi-view images across timesteps into a single grid-based representation; although this expands visual coverage, the compressed format can hinder fine-grained spatial-temporal understanding.
-Moreover, most existing methods output categorical labels without providing explicit reasoning traces, limiting interpretability and the ability to diagnose complex failure modes.
-Figure 1: Illustration of our Guardian model - a VLM fine-tuned on our constructed failure datasets. It detects planning failures (top) and execution failures (bottom) in robotic manipulation.
-To address these challenges, we introduce an automatic failure generation approach that produces diverse planning and execution failures, and we train a reasoning VLM Guardian on the resulting data for robust failure detection.
-Our method procedurally perturbs successful robot trajectories in both simulation and real-world settings to generate failures annotated with fine-grained categories and step-by-step reasoning traces.
-Using this pipeline, we construct three new datasets: RLBench-Fail (14K execution and 7K planning samples in the RLBench simulator [james2019rlbenchrobotlearningbenchmark]), BridgeDataV2-Fail (10K execution and 6K planning samples using the real-robot BridgeDataV2 data [walke2024bridgedatav2datasetrobot]), and UR5-Fail (570 execution and 370 planning samples), our collected real-robot failure dataset with policy-driven rollouts.
-We then develop Guardian, a reasoning VLM fine-tuned for failure detection.
-As illustrated in Fig.
-1, Guardian formulates failure detection as a visual question-answering task, reasoning over task instructions, subtask descriptions, and the visual scene.
-It leverages high-resolution, multi-view images to provide fine-grained and interpretable failure categorization.
-Guardian achieves state-of-the-art performance on both RLBench-Fail and BridgeDataV2-Fail, and generalizes in a zero-shot manner to RoboFail [liu2023reflect] and our UR5-Fail dataset.
-We further demonstrate Guardian’s plug-and-play capability as a feedback module by integrating it into a robotic manipulation system 3D-LOTUS++ [garcia2025generalizablevisionlanguageroboticmanipulation].
-Guardian significantly improves task success through planning and execution monitoring in both simulated and real-world tasks.
+Recent advances in Large Language Models (LLMs) [mistralsmall, grattafiori2024llama3herdmodels] and Vision-Language Models (VLMs) [zhu2025internvl3exploringadvancedtraining] have significantly improved vision-language robotic manipulation.
+Nevertheless, existing models remain vulnerable to diverse failures [sinha2023systemlevelviewoutofdistributiondata, Kawaharazuka_2024, kroemer2020reviewrobotlearningmanipulation] such as incorrect task decomposition, object confusion, or unstable grasps, which compound over long horizons and degrade real-world reliability.
+As a result, automatic failure detection and recovery has received growing research attention [liu2023reflect, chen2024automatingrobotfailurerecovery, duan2025aha, agia2024unpackingfailuremodesgenerative, etukuru2024robotutilitymodelsgeneral, ifailsense2026, zeng2025vifailback].
+Leveraging their strong generalization ability, LLMs and VLMs have been increasingly explored for failure detection.
+Some methods [liu2023reflect, etukuru2024robotutilitymodelsgeneral, duan2024manipulateanythingautomatingrealworldrobots], directly prompt pretrained foundation models to detect failures, optionally enhanced with chain-of-thought (CoT) reasoning [agia2024unpackingfailuremodesgenerative, nvidia2026cosmosreason2, cot2022] or multi-agent code generation [zhou2024code].
+While promising, these approaches suffer from a large domain gap: robotic observations differ substantially from web-scale pretraining data, and accurate failure detection requires fine-grained, embodied reasoning beyond generic visual understanding.
+Therefore, recent work [duan2025aha, ifailsense2026, robofac2025] has shifted toward fine-tuning VLMs on robot failure datasets to better bridge the gap.
+A fundamental bottleneck, however, is the scarcity of large-scale, high-quality failure data.
+Most robot learning datasets predominantly contain successful demonstrations [khazatsky2025droidlargescaleinthewildrobot, embodimentcollaboration2024openxembodimentroboticlearning, pumacay2024colosseumbenchmarkevaluatinggeneralization], providing limited failure examples.
+Collecting failures by rolling out policies is time-consuming and potentially unsafe, while manual curation [chen2024automatingrobotfailurerecovery, bu2025agibot] is labor-intensive and typically lacks diversity.
+Several prior approaches [duan2025aha, agia2024unpackingfailuremodesgenerative, dai2024racer] rely on simulated failure examples, but these suffer from sim-to-real gap [simtorealgapzhao2020], and provide limited coverage of both low-level execution errors and high-level planning failures [ifailsense2026].
+To address these limitations, we propose an automatic failure generation framework that synthesizes diverse planning and execution failures across simulated and real-world environments.
+Starting from successful demonstrations, we procedurally perturb task plans and subtask executions to create realistic failures, augmenting each example with structured step-by-step reasoning traces.
+This enables the construction of FailCoT, a large-scale failure reasoning dataset containing over 30K training examples.
+It includes RLBench-Fail built using the RLBench simulator [james2019rlbenchrobotlearningbenchmark] and BridgeDataV2-Fail derived from the BridgeDataV2 real-robot dataset [walke2024bridgedatav2datasetrobot], see Figure 1.
+FailCoT provides balanced success and failure samples, multi-view visual observations, and explicit CoT supervision for plan and subtask-level verification.
+Building on FailCoT, we develop Guardian, a multi-view reasoning VLM fine-tuned for unified planning and execution failure detection.
+Guardian formulates verification as a visual question answering problem: conditioned on task instructions, proposed plans or subtasks, and multi-view observations, it produces explicit reasoning traces to enhance predicting failures.
+To further support realistic evaluation, we introduce a new real-robot benchmark UR5-Fail constructed using the same failure generation framework.
+Guardian achieves state-of-the-art performance on three unseen real-world failure benchmarks, namely RoboFail [liu2023reflect], RoboVQA [sermanet2024] and UR5-Fail.
+When integrated as a plug-and-play verification module into a LLM-based manipulation system, Guardian improves task success in both simulation and real-robot experiments.
+Extensive ablations demonstrate the benefits of scaling structured, cross-environment failure reasoning data.
 In summary, our contributions are three-fold:
-  • We propose a novel robot failure generation approach that perturbs successful robot demonstrations in both simulation and real-world settings, and autonomously annotates failures with fine-grained categories and reasoning traces.
-This yields three new datasets for robot failure detection: RLBench-Fail, BridgeDataV2-Fail, and UR5-Fail.
-  • We introduce Guardian, a fine-tuned VLM for planning and execution failure detection.
-Guardian integrates high-resolution, multi-view visual inputs with explicit reasoning to predict fine-grained categorization.
-  • Guardian achieves state-of-the-art performance across four benchmarks, covering both in- and out-of-domain datasets.
-It further improves task completion of a vision-language manipulation framework in simulated and real robot tasks.
+  • We propose an automatic cross-environment failure synthesis framework that generates diverse planning and execution errors with structured reasoning supervision, resulting in the large-scale robot failure dataset FailCoT.
+  • We develop Guardian, a multi-view reasoning VLM fine-tuned on the FailCoT dataset for unified planning and execution failure detection.
+  • We show that scaling structured failure reasoning data yields state-of-the-art detection performance and improves task success when deployed as a plug-and-play verifier.
 We will release datasets, code, and models.
+Figure 1: Failure Data Generation Framework.
+We construct failure cases both online in simulation (RLBench), and offline on the real-world dataset (BridgeDataV2).
+For each positive example, given its correct plan and successful trajectory, we generate a corresponding incorrect plan and unsuccessful trajectory.
 II Related Work
+Vision-Language Robotic Manipulation.
+Recent advances in foundation models [zhu2025internvl3exploringadvancedtraining, openai2024gpt4ocard] have significantly improved vision–language robotic manipulation.
+End-to-end vision–language–action (VLA) policies such as Gr00T [groot2025] and $\pi_{0}$ [black2024pi] directly predict action sequences from 2D images and task instructions.
+To enhance spatial reasoning, 3D-based VLAs have further been proposed [garcia2025generalizablevisionlanguageroboticmanipulation, goyal2023rvtroboticviewtransformer].
+Notably, 3D-LOTUS++ [garcia2025generalizablevisionlanguageroboticmanipulation] achieves state-of-the-art performance on challenging generalizable manipulation tasks through a modular design that combines LLM-based task planning, visual grounding modules, and 3D-based execution policies.
+Despite this progress, robust robotic manipulation remains challenging, as planning mistakes and execution errors accumulate over long horizons [wu2025robomind].
+To improve robustness, recent approaches incorporate runtime monitoring of failures to trigger policy correction or retry [duan2024manipulateanythingautomatingrealworldrobots, etukuru2024robotutilitymodelsgeneral, dai2024racer].
+In this work, we advance automatic failure detection by scaling structured training data and demonstrate improved integration with manipulation policies.
+Robotic Failure Detection Methods.
+Early rule-based failure detection methods [de1998execution, gianni2011unified] struggle to generalize beyond predefined task structures.
+Learning-based approaches address this limitation and can be broadly categorized based on whether they require robot failure data for training.
+Training-free methods.
+One line of work performs out-of-distribution (OOD) detection [canwedetectfailurewithoutfailuredata2025] or temporal inconsistency detection [agia2024unpackingfailuremodesgenerative] over internal policy representations to flag failures without failure supervision.
+Another line prompts LLMs or VLMs for failure assessment [ahmad2025unifiedframework] using techniques such as hierarchical CoT reasoning [liu2023reflect, agia2024unpackingfailuremodesgenerative] or constraint-aware visual programming [zhou2024code].
+While these approaches avoid collecting robot failure data, they rely on policy-specific signals or prompt engineering, and cannot learn from robotic failure data for better accuracy.
+Training-based methods.
+Recent works [duan2025aha, ifailsense2026, nvidia2026cosmosreason2, gu2025safe, pmlr-v232-du23b, armor2026] fine-tune VLMs as failure detectors using annotated trajectories, but mostly target execution-stage verification.
+SuccessVQA [pmlr-v232-du23b] performs coarse task-level success assessment.
+AHA [duan2025aha] and I-Fail-Sense [ifailsense2026] compress multi-view inputs into a single concatenated image for detecting subtask-level execution failures or instruction–behavior misalignment.
+SAFE [gu2025safe] relies on internal policy representation to train a failure classifier.
+Cosmos-Reason [nvidia2026cosmosreason2] addresses general embodied reasoning trained with supervised fine-tuning and reinforcement learning, where failure detection is one downstream task.
+ARMOR [armor2026] uses multi-round self-refinement with separate detection and reasoning heads, but it trains only on post-execution failures and requires multiple inference passes per sample.
+Compared to prior methods, our Guardian model leverages large-scale failure reasoning data to enable multi-view, explicit reasoning for unified planning and execution verification, achieving state-of-the-art performance.
 Robot Failure Datasets.
-Existing robot failure datasets cover only limited settings.
-RoboFail [liu2023reflect] provides a small collection of hand-crafted simulation and real-world failures.
-Sentinel [agia2024unpackingfailuremodesgenerative] generates failures by rolling out a trained policy on out-of-distribution scenarios created through randomizing object scales and poses, but it includes only four tasks.
-AHA [duan2025aha] perturbs trajectories in RLBench simulator [james2019rlbenchrobotlearningbenchmark] to create over 49K image–query pairs, yet the data is restricted to simulation and does not include high-level planning failures.
-Moreover, none of these datasets provides step-by-step reasoning annotations, limiting the development of models capable of fine-grained failure reasoning.
-In this work, we propose an automated approach that generates large-scale datasets with diverse planning and execution failures, enriched with fine-grained categories and reasoning traces in both simulation and real-robot environments.
-Failure detection in robotic manipulation.
-Traditional monitoring methods rely on explicit models of tasks, identifying model deviations [de1998execution, gianni2011unified].
-Unlike LLM-based approaches, they depend on rigid model-based predictions.
-Recent works formulate failure detection as a Question Answering (QA) task, by zero-shot prompting LLMs and VLMs [agia2024unpackingfailuremodesgenerative, etukuru2024robotutilitymodelsgeneral, chen2024automatingrobotfailurerecovery, ma2024generative, shirai2024visionlanguageinterpreterrobottask, skreta2024replanroboticreplanningperception, mei2024replanvlmreplanningrobotictasks].
-For instance, REFLECT [liu2023reflect] summarizes multi-sensory inputs as texts, including audio, and feeds the text to an LLM, but its reliance on consecutive off-the-shelf models leads to accumulated errors.
-Consequently, efforts have been made to fine-tune VLMs specifically for robotic failure detection.
-SuccessVQA [pmlr-v232-du23b] fine-tunes Flamingo [flamingo] using both simulated and real-world data, but evaluates only the final task success, while Guardian is designed for more granular monitoring, checking the execution of each subtask.
-Closest to our work, AHA [duan2025aha] fine-tunes LLaVA-1.5 [liu2023visualinstructiontuning] on multi-class failures generated in RLBench.
-AHA concatenates multiple camera viewpoints and keysteps into a single image, leading to compressed visual inputs.
-Unlike AHA and SuccessVQA, which treat failure detection as direct classification, Sentinel [agia2024unpackingfailuremodesgenerative] and Cosmos-Reason1 [nvidia2025cosmosreason1physicalcommonsense] frame it as a reason-first process with explicit chain-of-thought (CoT) [10.5555/3600270.3602070] reasoning but focus only on single-view binary failure detection.
-Sentinel proposes a zero-shot CoT prompting framework to leverage off-the-shelf VLMs, while Cosmos-Reason1 is fine-tuned for physical reasoning.
-Extending prior approaches, Guardian unifies planning and execution verification via chain-of-thought reasoning over separated multi-view inputs, for multi-class failure detection.
-Learning-based robotic manipulation policies.
-Recent robotic policies have achieved remarkable performance on manipulation tasks [garcia2025generalizablevisionlanguageroboticmanipulation, embodimentcollaboration2024openxembodimentroboticlearning, ze20243ddiffusionpolicygeneralizable, black2025pi].
-However, their real-world applicability is still limited by planning and execution errors leading to task failures [wu2025robomind].
-Inspired by Manipulate-Anything [duan2024manipulateanythingautomatingrealworldrobots] and Robot Utility Models [etukuru2024robotutilitymodelsgeneral], which demonstrated significant benefits from integrating VLM verifiers into their pipelines, Guardian is designed as a plug-and-play verification module trained on specific failure data.
-It can seamlessly integrate into existing policies, enhancing robustness by verifying planning and execution stages.
-III Robot Failure Datasets Construction
-Figure 2: Failure Data Generation Pipeline.
-We introduce a novel generation pipeline generating failure cases both online in simulation (RLBench), and offline on the real-world dataset (BridgeDataV2).
-For each positive example, given its correct plan and successful trajectory, we generate a corresponding incorrect plan and unsuccessful trajectory.
+Collecting real-world robot failures at scale is challenging: policy rollouts are time-consuming, potentially unsafe, and require extensive manual annotation.
+RoboFail [liu2023reflect] provides a hand-crafted dataset spanning simulation and real settings, but covers limited tasks and failure modes.
+ViFailback [zeng2025vifailback] focuses on single-view, single-embodiment real-world diagnosis and requires substantial teleoperation effort.
+To reduce collection cost, several works rely on synthetic failure generation.
+Sentinel [agia2024unpackingfailuremodesgenerative] induces failures via out-of-distribution rollouts but covers only four tasks.
+SAFE [gu2025safe] utilizes the final sparse reward for policy rollouts and lacks dense failure supervision for each step.
+AHA [duan2025aha] perturbs trajectories in RLBench [james2019rlbenchrobotlearningbenchmark], generating large-scale purely simulated data, yet excludes high-level planning failures, and the dataset has not been publicly released.
+RoboFAC [robofac2025] adds reasoning annotations in simulation, with only a limited real-world subset via manual teleoperation.
+I-Fail-Sense [ifailsense2026] synthesizes failures from RLBench and DROID, but focuses primarily on semantic mismatches, leaving planning and low-level control errors underexplored.
+In contrast, we propose an automated pipeline that generates diverse planning and execution failures across both simulation and real robots, producing realistic failure modes at scale with multi-view observations and fine-grained, step-by-step reasoning supervision.
+III FailCoT: Cross-Environment Robot Failure Reasoning Datasets
 III-A Data Sources
@@
 Fig.
-2 (middle) illustrates successful episodes from the simulated and real robot datasets.
+1 (middle row) illustrates successful episodes from the simulated and real robot datasets.
 Simulated Data.
 We use the RLBench [james2019rlbenchrobotlearningbenchmark] simulator, selecting 52 tasks from RLBench-18Task [shridhar2022perceiveractormultitasktransformerrobotic] and GemBench [garcia2025generalizablevisionlanguageroboticmanipulation] benchmarks in our training data.
-For each task, we generate successful scripted trajectories with varied object placements and segment them into subtasks following 3D-LOTUS++ [garcia2025generalizablevisionlanguageroboticmanipulation].
+For each task, we generate successful scripted trajectories with varied object placements and segment them into subtasks following [garcia2025generalizablevisionlanguageroboticmanipulation].
 Real Robot Data.
@@
 A planning error denotes an incorrect decomposition of a task into subplans, whereas an execution error reflects unsuccessful completion of a subplan.
-Planning Failures. As shown in Fig. 2 (top), we construct five types of planning failures:
+Planning Failures. As shown in Fig. 1 (top row), we construct five types of planning failures:
   1. (1)
@@
 In simulation, we directly perturb subtask-level actions (Fig.
-2, bottom left), leveraging the simulator’s precise control.
+1, bottom left), leveraging the simulator’s precise control.
 A randomly selected subtask on the trajectory is modified using four failure modes:
@@
   4. (4)
-    Imprecise grasping/pushing – the gripper attempts to grasp or push the correct object by moving toward it and closing its jaws, but misses due to inaccurate positioning.
-Figure 3: Training examples. Guardian is trained on reasoning traces produced by our automatic data generation pipeline.
+    Imprecise grasping/pushing – the gripper attempts to grasp or push the correct object by moving toward it and closing its jaws, but misses it due to inaccurate positioning.
 For real robot data, modifying actions directly is impractical due to current limitations of image editing and generation models.
 Therefore, we perturb the subtask text instruction paired with the pre-recorded trajectory segment (Fig.
-2, bottom right) without direct robot control:
+1, bottom right) without direct robot control:
   1. (1)
@@
 Each execution example contains the task and subtask descriptions, plus pre-/post-action multi-view images.
+Figure 2: Planning and execution examples with CoT.
 III-C Chain-of-Thought (CoT) Generation
@@
 2 illustrates training examples with chain-of-thoughts verifying plan correctness and subtask completion.
-TABLE I: Robot failure datasets sizes across splits.
-Training Validation Test Dataset Exec Plan Exec Plan Exec Plan RoboFail [liu2023reflect] - - - - 153 30 RLBench-Fail 12358 5808 1000 500 1000 500 BridgeDataV2-Fail 7830 4880 1000 500 1000 500 UR5-Fail 400 200 30 30 140 140
 III-D Real-Robot, Policy-Driven Data Collection
-We curate UR5-Fail, a real-robot dataset, collected using a UR5 arm with three cameras.
-We run the 3D-LOTUS++ policy [garcia2025generalizablevisionlanguageroboticmanipulation] on 34 tasks, recording initial and final multi-view images for each subtask.
+To further support realistic evaluation, we curate UR5-Fail, a real-robot dataset, collected using a UR5 arm with three cameras.
+We run the 3D-LOTUS++ policy [garcia2025generalizablevisionlanguageroboticmanipulation] on 16 unique tasks, recording initial and final multi-view images for each subtask.
 Subtasks are manually labeled as success or failure to obtain execution failure data.
@@
 Unlike RoboFail [liu2023reflect], which is single-view and relies solely on teleoperation, UR5-Fail is three-view and features autonomous policy rollouts yielding more realistic failures.
-Figure 4: Failure mode distributions in real executions and our constructed data.
 III-E Dataset Statistics and Evaluation
-The resulting datasets, RLBench-Fail, BridgeDataV2-Fail, and UR5-Fail, contain balanced success/failure examples across both planning and execution, with fine-grained category labels and reasoning traces.
-Each dataset is split into training, validation, and test sets, with the validation and test sets featuring unseen tasks/environments to evaluate generalization.
-Dataset statistics are shown in Table I.
+FailCoT (RLBench-Fail, BridgeDataV2-Fail) and UR5-Fail, contain balanced success/failure examples across both planning and execution, with reasoning traces.
+FailCoT is split into training, validation, and test sets, with the validation and test sets featuring unseen tasks/environments to evaluate generalization, see Table I top.
+Table I bottom compares UR5-Fail with two existing real-world datasets and shows a more balanced distribution between execution and planning.
 To measure the quality and diversity of our synthetic datasets, i.e., whether the generated failures reflect real policy execution, we run the 3D-LOTUS++ policy [garcia2025generalizablevisionlanguageroboticmanipulation] on 92 RLBench tasks and manually annotate failure modes for 3 failure episodes per task.
 As shown in Fig.
-4, our designed failure modes reflect real failures, and the overall distribution of our synthetic and real failures remains similar.
-IV Method: The Guardian Model
-Figure 5: Left: Overview of the Guardian model architecture. Right: Integration of Guardian model into a robot manipulation pipeline for planning and execution verification.
+3, our designed failure modes reflect real failures, and the overall distribution of our synthetic and real failures remains similar.
+TABLE I: Statistics of the FailCoT dataset and three real-world robot failure detection benchmarks.
+Our constructed datasets (FailCoT and UR5-Fail) contain balanced success and failure cases covering both execution and planning errors.
+Training Validation Test
+Dataset Env. Exec Plan Exec Plan Exec Plan
+FailCoT (Ours)
+   RLBench-Fail Sim 12358 5808 1000 500 1000 500
+   BridgeDataV2-Fail Real 7830 4880 1000 500 1000 500
+Real-World Robot Failure Detection Benchmarks
+UR5-Fail (Ours) Real - - - - 140 140
+RoboFail [liu2023reflect] Real - - - - 153 30
+RoboVQA [sermanet2024] Real - - - - 357 -
+Figure 3: Failure mode distributions in real executions and our constructed data.
+IV Guardian: A multi-view reasoning VLM for robot failure detection
+Figure 4: Left: Overview of the Guardian model architecture.
+Right: Integration of Guardian model into a robot manipulation pipeline for planning and execution verification.
+The failure reasoning trace is reinjected to help with replanning.
 IV-A Problem Formulation
 We formulate robot failure detection as a visual question answering problem.
-For planning verification, given a high-level task instruction $T$, a proposed plan $P=(P_{1},\cdots,P_{N})$, and the initial visual context $I_{\text{start}}$, the model $\text{VLM}_{\text{plan}}$ must not only decide whether the plan is correct but also categorize the type of failure when it occurs:
-$\text{VLM}_{\text{plan}}(I_{\text{start}},T,P)\rightarrow(B_{\text{plan}},C_{\text{plan}}),$ (1)
-where $B_{\text{plan}}\in\{0,1\}$ indicates whether the plan is valid, and $C_{\text{plan}}$ specifies the failure category among the five planning failures constructed in Sec. III-B, shown in Fig. 4.
+For planning verification, given a high-level task instruction $T$, a proposed plan $P=(P_{1},\cdots,P_{N})$, and the initial visual context $I_{\text{start}}$, the model $\text{VLM}_{\text{plan}}$ must decide whether the plan is correct or not:
+$\text{VLM}_{\text{plan}}(I_{\text{start}},T,P)\rightarrow B_{\text{plan}}$ (1)
+where $B_{\text{plan}}\in\{0,1\}$ indicates planning success.
 For execution verification, given the task goal $T$, a subtask description $P_{i}$, and the visual observations before and after execution, $I_{\text{start}}$ and $I_{\text{end}}$, the model $\text{VLM}_{\text{exec}}$ similarly outputs
-$\text{VLM}_{\text{exec}}(I_{\text{start}},I_{\text{end}},T,P_{i})\rightarrow(B_{\text{exec}},C_{\text{exec}}),$ (2)
-where $B_{\text{exec}}\in\{0,1\}$ indicates execution success, and $C_{\text{exec}}$ denotes the execution failure category among the five execution failures constructed in Sec.
-III-B, shown in Fig.
-4.
+$\text{VLM}_{\text{exec}}(I_{\text{start}},I_{\text{end}},T,P_{i})\rightarrow B_{\text{exec}}$ (2)
+where $B_{\text{exec}}\in\{0,1\}$ indicates execution success.
 IV-B Model Architecture
-The Guardian model is built upon the state-of-the-art open-source VLM InternVL3-8B [zhu2025internvl3exploringadvancedtraining].
-As illustrated in Fig.
-5 (left), it consists of three key components: a text tokenizer that converts text prompts into discrete token embeddings, a visual encoder (InternViT-300M) that transforms individual images into visual embeddings, and a transformer-based LLM (Qwen2.5-7B) that processes the concatenated multimodal token sequence to predict the answer.
-Rather than concatenating multiple images into a single grid-based image as in AHA [duan2025aha], Guardian processes each image independently through the visual encoder.
-This design preserves fine-grained spatial details within each image and allows the model to explicitly reason about spatial and temporal changes for more accurate failure detection.
-Furthermore, unlike SuccessVQA [nvidia2025cosmosreason1physicalcommonsense] and AHA [duan2025aha] that output direct classifications, Guardian leverages an explicit reasoning trace before concluding success or failure.
+The Guardian model is built upon the open-source VLM InternVL3-8B [zhu2025internvl3exploringadvancedtraining].
+As shown in Fig.
+4 (left), it comprises three components: a text tokenizer that converts text into discrete token embeddings, a visual encoder (InternViT-300M) that transforms individual images into visual embeddings, and a transformer-based LLM (Qwen2.5-7B) that processes the concatenated multimodal tokens to predict the answer.
+Rather than concatenating multiple images into a single grid-based image as in prior work [duan2025aha, ifailsense2026], Guardian processes each image independently through the visual encoder.
+This design preserves fine-grained spatial details within each image and allows to explicitly reason about spatial and temporal changes for more accurate failure detection.
+Furthermore, instead of directly outputting classifications [duan2025aha, pmlr-v232-du23b, ifailsense2026], Guardian generates an explicit reasoning trace before concluding success or failure.
 IV-C Model Training
-We fine-tune Guardian on robot failure datasets using parameter-efficient Low-Rank Adaptation (LoRA) [hu2022lora], while freezing the visual encoder.
-Training minimizes cross-entropy loss for next-token prediction.
+We fine-tune Guardian on FailCoT using parameter-efficient Low-Rank Adaptation (LoRA) [hu2022lora], while freezing the visual encoder. Training minimizes cross-entropy loss for next-token prediction.
 Although CoT has shown promise to improve performance, it brings additional computation overhead.
 Inspired by prior work [chen2025trainingstrategiesefficientembodied], we explore three strategies for incorporating CoT into failure detection: (1) Vanilla: a baseline model trained and evaluated to directly predict final answers (A) without CoT; (2) Thinking: the model is trained and inferred with explicit reasoning, always generating CoT before A; (3) Dropout: in training, the model alternates between generating CoT+A and directly predicting A, while at test time, only A is produced.
+Results in Sec.
+V-C show that adding reasoning traces consistently improves performance.
+The Thinking strategy performs best but increases inference time, while Dropout offers a better speed-accuracy trade-off.
 IV-D Integration into Robotic Manipulation Framework
@@
 4 (right), Guardian can be inserted at each planning and subtask execution step to detect potential failures.
-Upon detection, it can trigger replanning or re-execute the corresponding motion policy to facilitate recovery.
+Upon detection, it can trigger replanning or re-execute the corresponding motion policy to facilitate recovery, and use its fine-grained failure reasoning as a hint to better replan.
 V Experiments
@@
 Evaluation datasets.
-We evaluate models on our newly constructed benchmarks RLBench-Fail, BridgeDataV2-Fail, and UR5-Fail, as well as the existing benchmark RoboFail [liu2023reflect], which is a manually created real-world failure dataset using a UR5 arm with a single camera.
-Evaluation metric. We report the mean binary detection accuracy for each dataset (success/failure) in Sec. V-B and discuss multi-class categorization in Sec. V-C.
+Our main evaluation focuses on three unseen real-world benchmarks: RoboFail [liu2023reflect], UR5-Fail, and RoboVQA [sermanet2024].
+RoboFail is a manually curated single-view UR5 failure dataset.
+UR5-Fail is our constructed multi-view real-robot dataset.
+RoboVQA (RVQA) is single-view and spans three embodiments: an Everyday Robots mobile manipulator, a human arm, and a human using a grasping tool.
+[^1] We use the RoboVQA test split restricted to execution success prediction.
+The original dataset also contains “planning” questions, but these focus on next-action/state prediction rather than plan verification.
+In ablations, we additionally report results on FailCoT testing splits.
+We use average classification accuracy as the metric.
 Implementation details.
-We trained our models using LoRA (rank 16, effective batch size 16) with the AdamW optimizer (weight decay 0.05), bf16 precision, and a cosine learning rate schedule peaking at $4\times 10^{-5}$.
-Training was conducted on the RLBench-Fail and BridgeDataV2-Fail sets (unless noted), completing in 5.5 hours on 4×H100 GPUs.
-During training, we randomly selected one or four views on RLBench-Fail data to mitigate overfitting to the multiple views.
-The best checkpoint is selected using the validation sets.
-TABLE II: Comparison of SOTA failure detection models.
-Execution and Planning binary accuracies are reported.
-denotes numbers from the AHA paper.
-Guardian is fine-tuned from InternVL3-8B on our constructed datasets, boosting the performance of the base model by 23%.
-(a) In-Domain datasets.
-Model RLBench-Fail BDV2-Fail Exec Plan Exec Plan Large-scale generalist models Qwen3-VL-235B-A22B 0.59 0.83 0.75 0.86 GPT4.1 0.63 0.87 0.72 0.86 Specialized / small models CLIP+MLP [pmlr-v139-radford21a] 0.65 0.53 0.58 0.54 Sentinel [agia2024unpackingfailuremodesgenerative] 0.57 - 0.57 - Cosmos-Reason1-7B [nvidia2025cosmosreason1physicalcommonsense] 0.54 0.60 0.53 0.61 InternVL3-8B 0.59 0.70 0.66 0.71 Guardian-8B-Thinking 0.83 0.87 0.85 0.91
-(b) Out-Of-Domain datasets.
-Model RoboFail [liu2023reflect] UR5-Fail Exec Plan Exec Plan Large-scale generalist models Qwen3-VL-235B-A22B 0.82 0.70 0.79 0.84 GPT4.1 0.82 0.67 0.79 0.88 Specialized / small models CLIP+MLP [pmlr-v139-radford21a] 0.42 0.43 0.51 0.51 AHA-13B [duan2025aha] 0.64 - - - Sentinel [agia2024unpackingfailuremodesgenerative] 0.79 - 0.74 - Cosmos-Reason1-7B [nvidia2025cosmosreason1physicalcommonsense] 0.67 0.70 0.58 0.63 InternVL3-8B 0.77 0.53 0.71 0.75 Guardian-8B-Thinking 0.86 0.70 0.77 0.89
+We fine-tune models using LoRA (rank 16, effective batch size 16) with AdamW (weight decay 0.05), bf16 precision, and a cosine schedule peaking at $4\times 10^{-5}$.
+Training is conducted on H100 GPUs using FailCoT unless otherwise specified.
+For RLBench-Fail, we randomly sample one or four views during training to mitigate view-specific overfitting.
+The best checkpoint is selected via validation accuracy.
+TABLE II: Comparison of failure detection models on unseen real-world benchmarks. Execution and planning accuracies are reported. denotes numbers from the original paper.
+Model Trained on RoboFail [liu2023reflect] UR5-Fail RVQA [sermanet2024]
+FailCoT Exec Plan Exec Plan Exec
+Closed-Source VLM
+GPT-4o 0.80 0.67 0.77 0.85 0.79
+GPT-4o +Sentinel-Video-QA [agia2024unpackingfailuremodesgenerative] 0.80 0.63 0.76 0.62 0.66
+Robotic Failure Detection VLMs
+RoboFAC-7B [robofac2025] 0.25 0.05 0.54 0.02 0.52
+AHA-13B [duan2025aha] 0.64 - - - -
+I-Fail-Sense-3B [ifailsense2026] 0.43 0.67 0.47 0.46 0.53
+Cosmos-Reason2-8B [nvidia2026cosmosreason2] 0.78 0.53 0.59 0.67 0.76
+CLIP+MLP [pmlr-v139-radford21a] 0.42 0.43 0.51 0.51 0.52
+I-Fail-Sense-3B [ifailsense2026] 0.76 0.52 0.55 0.6 0.58
+Cosmos-Reason2-8B [nvidia2026cosmosreason2] 0.82 0.70 0.65 0.83 0.77
+Guardian-8B 0.86 0.70 0.77 0.89 0.85
 V-B Comparison with State of the Art
 Compared methods.
-We compare our approach against SOTA general and specialized vision–language baselines.
-General baselines include small-scale (InternVL3-8B [zhu2025internvl3exploringadvancedtraining]), and large-scale models (Qwen3-VL-235B-A22B-Thinking [qwen3vl2025], GPT4.1 [openai2025gpt41]).
-Specialized baselines comprise Cosmos-Reason1-7B [nvidia2025cosmosreason1physicalcommonsense], AHA-13B [duan2025aha], Sentinel [agia2024unpackingfailuremodesgenerative], and an MLP baseline using CLIP text-image features [pmlr-v139-radford21a].
-We apply test-time CoT to the two large-scale models and Cosmos-Reason, as they are trained to produce reasoning traces and benefit from explicit reasoning at inference.
-Since AHA-13B is not publicly available, we report its binary accuracy on RoboFail from the original paper [duan2025aha].
-Sentinel is limited to execution monitoring and is not evaluated for planning.
+We compare against GPT-4o [openai2024gpt4ocard] and specialized robotic failure detectors including Cosmos-Reason2-8B [nvidia2026cosmosreason2], AHA-13B [duan2025aha], RoboFAC-7B [robofac2025], I-Fail-Sense-3B [ifailsense2026], Sentinel-Video-QA [agia2024unpackingfailuremodesgenerative], and CLIP+MLP [pmlr-v139-radford21a].
+AHA results come from the original paper, as the model is not publicly released.
+For the other methods, we run their released checkpoints or train models with the released codebase.
 Results.
-Table II reports in-domain (ID) and out-of-domain (OOD) binary accuracies.
-Table II(a) shows that our Guardian model, fine-tuned from InternVL3-8B on failure datasets, surpasses/matches the larger Qwen3VL and GPT4.1 on in-domain execution and planning sets.
-In Table II(b) we evaluate models zero-shot on OOD datasets: Guardian attains 0.86/0.70 (Exec/Plan) binary accuracy on RoboFail and 0.77/0.89 on UR5-Fail, outperforming/matching larger general-purpose and specialized VLMs.
-Sentinel, applied atop Qwen3-VL, underperforms its base VLM, likely due to Sentinel’s rigid CoT prompt.
-Cosmos-Reason, trained on single-view binary failures, often overlooks fine-grained manipulator/object state changes.
-AHA, solely tuned to simulation failure data, does not generalize well to the real robot failures of RoboFail.
+Table II reports performance on the test sets.
+GPT-4o achieves strong performance due to its scale and general reasoning ability.
+However, applying the Sentinel-Video-QA [agia2024unpackingfailuremodesgenerative] self-interrogation prompting degrades accuracy as it constrains the reasoning ability of the original model.
+Models trained exclusively on simulated failures (RoboFAC [robofac2025] and AHA [duan2025aha]) show limited transfer to real-robot benchmarks.
+Both rely on simulation-only perturbations, which likely restrict generalization to unseen real-world manipulators and sensor noise.
+I-Fail-Sense [ifailsense2026] is trained on both simulation and real-world trajectories, but its supervision focuses primarily on semantic misalignment detection rather than structured planning or low-level control failures, which likely limits its performance on the benchmarks.
+Cosmos-Reason [nvidia2026cosmosreason2] performs competitively, reflecting strong physical reasoning capabilities, but it is optimized for broad embodied reasoning rather than only for failure verification.
+Since prior failure detection models are trained on different datasets, we further fine-tune representative open-source models on the same FailCoT dataset to isolate the effects of data and architecture (Table II).
+We also include a lightweight CLIP+MLP baseline, which performs substantially worse, highlighting the necessity of large vision–language models.
+Notably, training on FailCoT consistently improves all methods, underscoring the importance of well-curated, cross-environment data with broad failure coverage.
+Guardian achieves the strongest overall performance across RoboFail, UR5-Fail, and RoboVQA.
+Compared to I-Fail-Sense [ifailsense2026], Guardian preserves multi-view spatial structure and produces explicit chain-of-thought reasoning, enabling structured subtask-level verification.
+Compared to Cosmos-Reason, which is pretrained for embodied reasoning and primarily developed in single-view settings, Guardian leverages an InternVL backbone with explicit multi-view supervision, which likely explains its stronger fine-grained failure detection performance even under identical training data.
 TABLE III: Impact of the Guardian fine-tuning data mix on the binary accuracy averaged over planning and execution.
-Training Data RLBench BDV2 Robo UR5 RLBench BDV2 UR5 -Fail -Fail -Fail -Fail 0.65 0.69 0.65 0.73 0.80 0.72 0.69 0.74 0.66 0.87 0.72 0.69 0.85 0.88 0.78 0.83 0.85 0.88 0.74 0.85
-Figure 6: Training data size impact on the Execution binary accuracy across benchmarks.
-Figure 7: Confusion matrices for Guardian’s failure categorization for plans and executions. Values are percentages; rows denote ground-truth classes and columns the predicted classes.
+Training Data RLBench BDV2 Robo UR5 Robo RLBench BDV2 -Fail -Fail -Fail -Fail -VQA 0.65 0.69 0.65 0.73 0.75 0.82 0.70 0.69 0.72 0.66 0.65 0.86 0.71 0.68 0.77 0.85 0.88 0.78 0.83 0.85
+Figure 5: FailCoT training data size impact on the binary accuracy averaged over planning and execution.
 V-C Failure Data Ablations
-In this section, we examine the composition and scale of our data, and discuss the multi-class categorization.
-Training data mixture.
-Table III shows the impact of fine-tuning datasets.
-Row 1 is the baseline InternVL 3-8B without fine-tuning.
-Fine-tuning on simulated RLBench-Fail data (row 2) significantly improves the performance on the same dataset, but the gains do not transfer well to the other real-world datasets due to the domain gap.
-A similar trend is observed when training solely on BridgeDataV2-Fail in row 3.
-In contrast, combining both datasets during fine-tuning in the fourth row leads to significant gains in domain and improves generalization to the OOD benchmarks.
-This indicates the importance of generating diverse planning and execution data.
-Further fine-tuning with the small real-robot UR5-Fail dataset improves performance on UR5-Fail while maintaining comparable accuracy on the other datasets.
-Training data size.
+We next analyze how training data composition (simulation and real data), and dataset scale influence cross-environment generalization while keeping the architecture fixed.
+Data composition.
+Table III compares InternVL3-8B without fine-tuning, with single-source training, and with the full FailCoT dataset.
+Without fine-tuning, performance is moderate.
+Training only on RLBench-Fail improves simulated results but transfers weakly to real-robot datasets.
+Training only on BridgeDataV2-Fail (BDV2) improves real-world performance but shows limited simulation transfer.
+Combining both datasets yields consistent gains across RoboFail, UR5-Fail, and RoboVQA.
+The same trend holds for other architectures in Table II, emphasizing the importance of cross-environment composition and broad coverage of failure supervision.
+Dataset scaling.
 Fig.
-6 presents the scaling behaviors of the training data size on the execution binary accuracies across RLBench-Fail, BridgeDataV2-Fail, RoboFail, and UR5-Fail.
-We observe a scaling behavior on in-domain and out-of-domain sets when fine-tuning on our procedurally generated data.
-This suggests that scaling the generated data size further could improve model performance.
-Multi-class Categorization.
-Since our datasets contain fine-grained failure categorization, Guardian supports multi-class failure prediction.
-The confusion matrices in Fig.
-7 show strong overall discrimination across categories.
-As shown in the Fig.
-7(left), most execution categories are identified with high accuracy.
-The main weakness lies in the VLM’s visual grounding of subtle cues: no gripper close is confused with imprecise grasping/pushing, and wrong object state or placement is often mistaken for success.
-For planning shown in Fig.
-7(right), wrong object state or placement, wrong order, missing subtask, and wrong object manipulated show higher confusion, particularly in long-horizon plans.
-Improving the model’s commonsense and temporal reasoning would further enhance its ability to capture these nuanced dependencies.
-TABLE IV: Image representation impact on the Execution binary accuracy (number of views and image format). The AHA method [duan2025aha] equals row 3 (4-view concatenation).
-Views
-Image
-Format
-RLBench-Fail
-Execution
-1 concat 0.61
-separated 0.63
-4 concat (AHA [duan2025aha]) 0.72
-separated 0.81
-Figure 8: Verification with Guardian during online task execution on RLBench. Left: Successful correction of a generated plan. Right: Successful correction of a subtask execution.
-V-D Failure Detection Method Ablations
-In this section, we compare visual representation choices and reasoning strategies for VLM-based failure detection.
-Multi-view images.
-Table IV investigates the image representation choice on failure detection.
-We fine-tune and evaluate InternVL3-8B solely on RLBench-Fail execution data.
-We vary the number of viewpoints (one or four) and the multi-image format, either separated as in Guardian, or concatenated into a single image as in AHA [duan2025aha].
-For concatenated inputs, we represent the views as rows, and start/end images as columns.
-Using multi-view images consistently outperforms their single-view counterpart.
-In the single-view setting, separating or concatenating two images (start and end) yields comparable performance, as the resolution loss from concatenation is minimal.
-However, in the four-view setting, separating the images significantly boosts the performance by 9 points over concatenation.
-Concatenating images not only shifts from natural image distributions but also compresses visual information in limited image resolution.
-TABLE V: Comparison of training and test-time strategies with and without CoT (Sec.
+5 demonstrates consistent scaling behavior as the amount of FailCoT data increases.
+Performance on in-domain and unseen real-world datasets improves steadily with dataset size.
+This indicates that scaling structured failure generation remains a promising direction for improving cross-environment generalization.
+TABLE IV: Comparison of training and test-time strategies with and without CoT (Sec.
 IV-C).
 “A” denotes the final answer.
-We report the average binary accuracy and inference time [seconds/sample] on one H100.
+We report the binary accuracy and inference time (seconds/sample) on one H100 averaged across all test sets.